OpenStudy (anonymous):

int \frac{ e ^{3x} }{(1+e ^{3x} )^{8}} dx

OpenStudy (loser66):

let \[u =1+e^{3x}\rightarrow du =3e^{3x}dx\] and solve as usual

OpenStudy (anonymous):

ok, thats where i got

OpenStudy (anonymous):

is it 1/u^8 ?

OpenStudy (loser66):

\[\int\frac{e^3x}{(1+e^{3x})^8}dx\]is it yours?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, so i was thinking right....

OpenStudy (loser66):

yes, it turns \[\int \frac{1}{3}u^{-8}du\]

OpenStudy (anonymous):

oh ok,...thats what i didnt see

OpenStudy (loser66):

so, you can step up from now, right?

OpenStudy (anonymous):

yes, just substitute and that is final answer, right?

OpenStudy (anonymous):

u can go to denominator

OpenStudy (loser66):

take integral of this, it s\[\frac{1}{3}\frac {u^{-7}}{(-7)} + C\\=\frac{-1}{21u^7}+C\\=\frac{-1}{21(1+e^{3x})^7}\]

OpenStudy (loser66):

+ C at last

OpenStudy (anonymous):

great. your awsome. thanks

OpenStudy (loser66):

yw